OpenStudy (anonymous):

Determine if the following is always, sometimes, or never true. y = -1/5 x + 3 is a function. explain please

OpenStudy (calculusfunctions):

First tell me what you think it is.

OpenStudy (anonymous):

yes

OpenStudy (calculusfunctions):

@StupidGenius can we please teach and not just give away the answers?

OpenStudy (anonymous):

well i think im suppose to graph it and do the vertical line test but i dont know how to graph equations:(

OpenStudy (calculusfunctions):

Let me give you a hint. You don't need to graph. Hint: If the exponent of y is 1, then it is always a function, except when it's y = ±√x.

OpenStudy (calculusfunctions):

I can give write you a proper definition of a function which you may find complicated at the level you are right now (I'm assuming you're in Gr. 9 or 10) but if you understand the hint, then you should be fine.

OpenStudy (anonymous):

um there is not exponent

OpenStudy (calculusfunctions):

When you don't see an exponent, the exponent is an "invisible" 1.

OpenStudy (anonymous):

oooooh so its a function

OpenStudy (calculusfunctions):

For example y means y to the exponent 1. Similarly x means x to the exponent 1. But in order for the equation to represent a function, only the exponent of y has to be 1. x can have any exponent. Understand?

OpenStudy (anonymous):

wait so as long as its y=such and such then its a funcyion but if its y^7=such and such its nit a function?

OpenStudy (anonymous):

kinda confused

OpenStudy (calculusfunctions):

You've got it! See now you've really earned a medal!

OpenStudy (calculusfunctions):

Why are you all of a sudden confused? I just gave you a medal and all! Now I'm confused!

OpenStudy (calculusfunctions):

I'm confused as to why you're confused.

OpenStudy (anonymous):

but all my problems include y=such and such and not all of them are functions

OpenStudy (calculusfunctions):

Go ahead and post them.

OpenStudy (calculusfunctions):

I told you that y = ±√x is an exception.

OpenStudy (anonymous):

i know this is gonna sound stupid but that is the onlhy ine i can find now

OpenStudy (calculusfunctions):

OK well then you need to trust me because FYI I am a teacher.

OpenStudy (calculusfunctions):

I wouldn't mislead you.

OpenStudy (anonymous):

alrighty thanks for the help:)

OpenStudy (calculusfunctions):

No problem! Do you need anything else right now?

OpenStudy (anonymous):

let me see

OpenStudy (anonymous):

waity so the answer is always?

OpenStudy (calculusfunctions):

YES!

OpenStudy (anonymous):

alright What is the slope of the graph for the equation y = -5x + 8?

OpenStudy (calculusfunctions):

y = mx + b is the equation of a line with slope m and y-intercept b. Knowing this, what do you think the slope is for the equation of y = -5x + 8

OpenStudy (anonymous):

-5..oh wait haha

OpenStudy (calculusfunctions):

YES! STOP DOUBTING YOURSELF AND HAVE MORE FAITH IN YOUR ABILITIES!!!!!!

OpenStudy (anonymous):

its that i have had such a hard time with math for a such a long time..also im really tired.

OpenStudy (calculusfunctions):

No worries! Just believe in yourself!

OpenStudy (anonymous):

question what is the equation for point slope form?

OpenStudy (anonymous):

thanks:)

OpenStudy (calculusfunctions):

\[y -y _{1}=m(x -x _{1})\] is the slope-point form of the equation of a line where m is the slope of the line and\[(x _{1}, y _{1})\]is a point on the line.

OpenStudy (anonymous):

oh ok so its like m=y-y1/x-x1

OpenStudy (calculusfunctions):

PERFECT!!!!! See! I wish the system allowed me to award you a second medal!

OpenStudy (anonymous):

haha thanks:)

OpenStudy (calculusfunctions):

I hope you're feeling more confident!

OpenStudy (anonymous):

i am thanks for all you help i have just one more question is the slope of a horizonal line 0?

OpenStudy (calculusfunctions):

Yes!

OpenStudy (calculusfunctions):

Then what is the slope of a vertical line?

OpenStudy (anonymous):

undefined

OpenStudy (anonymous):

can only be on my comp till9:30 cuss it has a block thing so i gotta get off but thanjk you so much for all your help:)i really appreciate it!!

OpenStudy (calculusfunctions):

WOW!!! You're smarter then you're letting on @EmmaH Am I being punked? LOL

OpenStudy (anonymous):

no i just do research:)

OpenStudy (anonymous):

and get help from nice people

OpenStudy (anonymous):

got to go bye :)

OpenStudy (calculusfunctions):

No problem!! You're welcome! Let me know if you need anything else! Bye!